A student measured a mass of a metal bar as 57.21
Answer #1

4. Density = mass / volume

Volume = water displaced by metal when placed in water filled graduated cylinder

= 53.70 - 48.60 = 5.1 ml

Now, calculate density, d = 57.21 / 5.1 = 11.22 g/ml

Density of metal is near to that of lead. Hence we can say given metal is lead.

5. Volume = 9.50 ml

We need value of density of rubbing alcohol to calculate the mass

Density of rubbing alcohol = 0.7854 g/ml

Now, d = m / v

Mass = density x volume = 0.7854 x 9.50 = 7.46 g.